# Flow-Enabled Generalization to Human Demonstrations in Few-Shot Imitation Learning

### v1

@@ Abstract
11footnotetext: The Australian National University. Correspondence to runze.tang@anu.edu.au
@@ I INTRODUCTION
As a fundamental subclass of Imitation Learning (IL), behavior cloning (BC) is a widely adopted strategy for offline policy learning from demonstrations [56]. While BC can address tasks with complex dynamics, such as deformable object manipulation, it usually requires dozens to hundreds of demonstrations for simple tasks [46, 10, 7], and hundreds of thousands to achieve robust generalization across diverse environments [6, 65]. However, collecting large-scale datasets, even via crowdsourcing, is often cost-prohibitive given the specialized equipment for each demonstrator [6, 65, 17]. Therefore, many studies focus on using videos of human manipulation as a substitute for robot demonstrations [3, 39, 55].
With 3D spatial information, point cloud-based methods generally have better data efficiency and generalization capabilities than image-based methods [64]. The image convolution features of the robot and human hand regions are inevitably different. Therefore, distribution adaptation methods are required to align image representations [45, 39]. While in a point cloud, the robot or human hand is represented by points in the air. Combined with segmentation methods, point clouds become a well-suited representation for cross-embodiment learning [55]. Flow, that is, the trajectories of points, has been confirmed to be an effective representation that describes motions to bridge human videos and robot demonstrations [39, 55, 53, 48, 63]. However, most prior work has focused on predicting flow on the object or robot arm solely to achieve cross-embodiment learning. Focusing solely on the object’s motion often overlooks the robot’s pre-grasp motion [55]. Conversely, only considering the motion of the robot omits the details of the interaction with the object [39]. In this work, we present a point cloud-based method that enables flow prediction for any points in the scene and cross-embodiment learning.

### v2

# arXiv:2602.10594v2[cs.RO] 27 Feb 2026

@@ Abstract
11footnotetext: The Australian National University {runze.tang, penny.kyburz}@anu.edu.au
@@ I INTRODUCTION
As a fundamental subclass of Imitation Learning (IL), behavior cloning (BC) is a widely adopted strategy for offline policy learning from demonstrations [56]. While BC can address tasks with complex dynamics, it usually requires dozens of demonstrations for simple tasks [46, 10, 7], and thousands to achieve robust generalization [6, 65]. However, collecting large-scale datasets is often cost-prohibitive given the specialized equipment for each demonstrator [6, 65, 17]. Therefore, many studies focus on using videos of human manipulation as a substitute for robot demonstrations [3, 39, 55].
With 3D spatial information, point cloud-based methods generally have better data efficiency and generalization capabilities than image-based methods [64]. The image convolution features of the robot and human hand regions are inevitably different. Therefore, distribution adaptation methods are required to align image representations [45, 39]. While in a point cloud, the robot or human hand is represented by points in the air. Combined with segmentation methods, point clouds become a well-suited representation for cross-embodiment learning [55]. Flow, that is, the trajectories of points, has been confirmed to be an effective representation that describes motions to bridge human and robot demonstrations [39, 55, 53, 48, 63]. However, most prior work has focused on flow on the object or robot arm solely to achieve cross-embodiment learning. Focusing solely on the object’s motion often overlooks the robot’s pre-grasp motion [55]. Conversely, only considering the motion of the robot omits the details of the interaction with the object [39]. In this work, we present a point cloud-based method that enables any-point flow prediction and cross-embodiment learning.
@@ VI Conclusion
ACKNOWLEDGMENT
We would like to thank the Robotics@ANUComputing group for providing the experimental equipment and workspace. We also thank Hanna Kurniawati, Miaomiao Liu, and Rahul Shome for their helpful feedback and valuable discussions.